Jefferson Lab HALL A Logbook for October 02 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 89387 10/31/02 16:59 Eric Energy Lock A/C
  • 89383 10/31/02 16:00 LeRose Shift Summary
  • 89376 10/31/02 12:37 LeRose 1C12 OTR
  • 89373 10/31/02 12:17 LeRose FPP Thresholds off/on
  • 89371 10/31/02 10:49 LeRose Replaced ep turbo
  • 89370 10/31/02 09:33 LeRose Compton Vacuum
  • 89366 10/31/02 08:04 H. Ibrahim Owl Shift Summary
  • 89360 10/31/02 05:39 H. Ibrahim Magnet Settings for Q3D_f40l Kinematics
  • 89348 10/31/02 01:18 pawel Run 2782
  • 89347 10/31/02 01:10 atrig Trigger-Download
  • 89338 10/31/02 00:02 jpchen Swin shift summary
  • Wednesday

  • 89327 10/30/02 18:22 paschke/kauf physics replay kumac corrected
  • 89321 10/30/02 16:03 Boeglin Shift Summary
  • 89312 10/30/02 11:48 Reitz ep energy measurement
  • 89303 10/30/02 08:06 H. Ibrahim Owl Shift Summary
  • 89297 10/30/02 05:52 peter target IOC reboot
  • 89281 10/30/02 00:04 paschke swing shift summary
  • Tuesday

  • 89274 10/29/02 18:55 paschke Non-appearance of end-of-run log
  • 89271 10/29/02 16:03 Higinbotham Shift Summary
  • 89265 10/29/02 11:49 M Jones moved tc_r to slot 16 chan0 roc 2
  • 89260 10/29/02 08:03 H. Ibrahim Owl Shift Summary
  • 89248 10/29/02 04:40 H. Ibrahim Magnet settings for Q3D_f30r kinematics
  • 89247 10/29/02 04:37 atrig Trigger-Download
  • 89235 10/29/02 00:55 H. Ibrahim Magnet settings for Q3D_f20r kinematics
  • 89234 10/29/02 00:52 H. Ibrahim Beam is not stable.
  • 89232 10/29/02 00:25 atrig Trigger-Download
  • 89230 10/29/02 00:04 weinstein magnet and beam problems
  • Monday

  • 89225 10/28/02 21:21 weinstein magnet ramp problem solved
  • 89223 10/28/02 18:40 weinstein magnet crashes
  • 89222 10/28/02 17:41 M Jones modified pk/e01020_phys_paw.kumac
  • 89221 10/28/02 17:24 Higinbotham Shift Summary
  • 89219 10/28/02 16:09 M Jones moved tc_r from slot 17 on roc1 to slot 16
  • 89214 10/28/02 15:02 Higinbotham Dipole Restored
  • 89209 10/28/02 14:08 bogdanw right arm VDC
  • 89207 10/28/02 12:14 Higinbotham Left Magnet Down
  • 89202 10/28/02 11:30 Adaq number of events problem
  • 89201 10/28/02 11:01 M Jones problem with run 2812 ( flaky ROC2?)
  • 89196 10/28/02 09:24 Higinbotham Right Wire Chamber
  • 89193 10/28/02 08:25 H. Ibrahim Kinematics corrections
  • 89192 10/28/02 08:13 Eric Magnet screen for Q3D_f20r
  • 89191 10/28/02 08:11 H. Ibrahim Owl Shift Summary
  • 89186 10/28/02 06:41 atrig Trigger-Download
  • 89185 10/28/02 06:37 a-onl Trigger-Download
  • 89183 10/28/02 05:37 H. Ibrahim Epics Time Chart
  • 89182 10/28/02 05:17 Eric Magnet screen for Q3D_f20l
  • 89169 10/28/02 00:17 Ulmer Swing Shift Summary
  • Sunday

  • 89167 10/27/02 23:36 johna spec_r_tc strange behavior
  • 89162 10/27/02 21:27 Ulmer Kin Q3D_f10l mislabeled
  • 89154 10/27/02 21:00 atrig Trigger-Download
  • 89142 10/27/02 18:42 Ulmer 1H(e,e'p) Elastic cross sections
  • 89141 10/27/02 18:40 suleiman parity target test
  • 89137 10/27/02 17:27 atrig Trigger-Download
  • 89135 10/27/02 16:58 I. QATTAN Changing from 4cm LH2 to 15 cm LD2 and JTV settings.
  • 89130 10/27/02 16:12 S. Frullani day shift summary
  • 89120 10/27/02 11:25 reinhold move target from 15cm H2 to 4cm H2
  • 89111 10/27/02 08:12 M Jones owl shift summary
  • 89103 10/27/02 06:38 Markowitz Loop 1 ready for beam
  • 89102 10/27/02 06:24 Markowitz further target motion addendum
  • 89101 10/27/02 06:12 Markowitz Target IOC reboot
  • 89100 10/27/02 05:53 Markowitz change to 15 cm LH2
  • 89092 10/27/02 03:49 Markowitz target change to 4 cm LH2
  • 89090 10/27/02 03:43 M Jones harp scan
  • 89085 10/27/02 02:23 Markowitz target move to 15 cm LD2
  • 89076 10/27/02 01:55 M JOnes rebooted iocha48 for Q1 Left and right arm
  • 89079 10/27/02 01:36 atrig Trigger-Download
  • 89078 10/27/02 01:14 Markowitz change to optics target
  • 89070 10/27/02 00:29 johna Swing shift summary
  • Saturday

  • 89067 10/26/02 22:46 Punjabi magnet settings
  • 89057 10/26/02 19:03 johna Beam is back
  • 89053 10/26/02 16:04 Eric Voutier Shift summary
  • 89041 10/26/02 08:10 R. Michaels updated "spot++"
  • 89040 10/26/02 08:05 M Jones owl shift summary
  • Friday

  • 89022 10/25/02 23:50 reinhold set target JT2 and Fan2
  • 89021 10/25/02 23:47 reinhold moved target to 15cm D2
  • 89020 10/25/02 23:46 M Jones software failed in harp scan
  • 89019 10/25/02 22:46 reinhold MCC would like to sell beam,
  • 89018 10/25/02 18:15 reinhold Q1 up again
  • 89017 10/25/02 17:10 reinhold Q1 won't come back
  • 89016 10/25/02 16:37 reinhold Q1 trip
  • 89015 10/25/02 16:05 paschke Shift Summary
  • 89013 10/25/02 11:40 paschke Q1 Left won't come up
  • 89012 10/25/02 10:17 paschke Q1 trip: what to do
  • 89011 10/25/02 08:59 paschke Target put "to sleep"
  • 89009 10/25/02 08:18 M Jones owl shift summary
  • 89005 10/25/02 06:58 M JOnes Turned off beamline hv card 1 ch 0 and 1
  • 89002 10/25/02 06:53 Markowitz target move to 15 cm LD2
  • 89000 10/25/02 06:25 Markowitz target move to optics
  • 88999 10/25/02 06:24 atrig Trigger-Download
  • 88994 10/25/02 05:00 M Jones FFB on
  • 88992 10/25/02 04:18 M Jones FFB off
  • 88985 10/25/02 00:42 M Jones OTR at H04 in
  • 88982 10/25/02 00:06 johna Swing shift summary
  • Thursday

  • 88979 10/24/02 23:41 reinhold target fan warning
  • 88978 10/24/02 23:37 reinhold moved target to 15cm D2
  • 88973 10/24/02 23:14 atrig Trigger-Download
  • 88972 10/24/02 22:59 reinhold moved target to optics
  • 88969 10/24/02 22:51 johna beamline HV trip during run 2736
  • 88964 10/24/02 21:48 atrig Trigger-Download
  • 88956 10/24/02 17:50 reinhold MCC announces G0 dump test
  • 88953 10/24/02 17:14 ulmer VDC spikes - where are they?
  • 88952 10/24/02 16:21 suleiman More LD2 parity test
  • 88951 10/24/02 16:07 LeRose Shift summary
  • 88948 10/24/02 14:19 Saha Hall A Arc Energy Measurement
  • 88946 10/24/02 11:18 jpchen Cryotarget fan problem fixed by target group, back to 60 Hz
  • 88944 10/24/02 10:38 LeRose Q1 reset
  • 88940 10/24/02 08:53 paschke Q1 trip: what to do.
  • 88938 10/24/02 08:02 peter shift summary
  • 88936 10/24/02 07:02 choi Left Q1 tripped
  • 88934 10/24/02 06:32 choi HV alarm
  • 88915 10/24/02 00:16 Eric Voutier Shift summary
  • Wednesday

  • 88912 10/23/02 23:09 A. Camsonne Target moved to 15 cm LD2
  • 88909 10/23/02 22:50 atrig Trigger-Download
  • 88908 10/23/02 22:45 A. Cams Target moved to optics target
  • 88907 10/23/02 22:43 reitz ep energy measurement failed
  • 88906 10/23/02 22:29 Eric Magnet screen Q3_d50
  • 88905 10/23/02 21:54 A. Camsonne Target moved to empty target
  • 88904 10/23/02 21:26 A. Camsonne Target movement failure
  • 88894 10/23/02 18:28 A. Camsonne Target IOC reboot
  • 88889 10/23/02 17:43 M Jones turned off FPP HV
  • 88887 10/23/02 16:19 M Jones modified ek/e01020_det.kumac
  • 88886 10/23/02 16:03 johna Shift summary
  • 88881 10/23/02 14:23 ulmer Right-arm VDC wire spectra
  • 88878 10/23/02 14:13 johna Gas alarm (out of nitrogen)
  • 88877 10/23/02 13:27 johna MCC turns on energy lock and FFB
  • 88874 10/23/02 12:59 johna FFB was turned off at 8:00 this morning.
  • 88873 10/23/02 12:43 johna Beam steering during run 2705
  • 88870 10/23/02 11:48 paschke Controlled Access summary
  • 88869 10/23/02 11:10 M Jones update dhist detmap
  • 88865 10/23/02 10:31 johna Tweaked position of left arm collimator
  • 88856 10/23/02 08:03 peter shift summary
  • 88839 10/23/02 00:41 M Jones look at tc_l versus tc_r
  • 88837 10/23/02 00:09 R. Michaels adaqfs cleanup
  • 88836 10/23/02 00:09 Eric Voutier Shift Summary
  • Tuesday

  • 88833 10/22/02 23:20 A. Camsonne Target moved to 15 cm LD2
  • 88830 10/22/02 23:02 atrig Trigger-Download
  • 88829 10/22/02 22:59 a-onl Trigger-Download
  • 88828 10/22/02 22:58 Eric Q3_d40 Magnet screen
  • 88827 10/22/02 22:47 R. Michaels adaqfs problems
  • 88824 10/22/02 22:23 A. Camsonne Target moved to optics
  • 88811 10/22/02 18:49 suleiman Parity Boiling Test
  • 88804 10/22/02 16:26 M Jones added variable edtm_patt to espace
  • 88802 10/22/02 16:01 ulmer Day Shift Summary
  • 88801 10/22/02 16:00 gilman LD2 beam compensation
  • 88800 10/22/02 15:54 gilman LD2 cryotarget
  • 88799 10/22/02 15:49 Ulmer Right arm scint. hits
  • 88797 10/22/02 15:43 gilman cryotarget status
  • 88796 10/22/02 15:36 gilman cryotarget beam compensation
  • 88794 10/22/02 15:27 M Jones modified espace to have variable for new trig pattern
  • 88791 10/22/02 15:07 atrig Trigger-Download
  • 88790 10/22/02 14:47 Saha Magnet Tools screen for Run 2676
  • 88788 10/22/02 14:39 Saha SPOT++ for Optics Run 2676
  • 88787 10/22/02 14:39 Saha Harp Scan for Optics Run 2676
  • 88784 10/22/02 14:06 Saha Spot++ for Optics Run 2675
  • 88783 10/22/02 14:06 Saha Harp Scan for Optics Run 2675
  • 88782 10/22/02 14:01 Saha Magnet Tools for Run 2675 - Optics Target Run
  • 88779 10/22/02 13:55 Saha Harp Scan
  • 88778 10/22/02 12:34 gilman cryotarget followup
  • 88777 10/22/02 12:30 R. Michaels raw triggers plugged into 1877
  • 88776 10/22/02 12:24 gilman and j cryotarget - turning on deu loop
  • 88774 10/22/02 10:55 gilman cryotarget
  • 88773 10/22/02 10:24 Ulmer Harp # 6 fixed?
  • 88772 10/22/02 08:47 ulmer Harp Scan
  • 88770 10/22/02 08:11 peter shift summary
  • 88747 10/22/02 05:17 choi Strange strip in SPOT
  • 88735 10/22/02 02:20 Eric LD2 Solid ??
  • 88734 10/22/02 02:16 Eric Beam (UN)stability
  • 88729 10/22/02 01:50 Choi Vacuum problem prevents high current for the moment
  • 88712 10/22/02 00:21 Eric Voutier Shift Summary
  • Monday

  • 88700 10/21/02 23:11 Suleiman Parity Test is Done
  • 88699 10/21/02 23:09 Suleiman Parity Target Boiling Test
  • 88696 10/21/02 22:41 Eric LH2 Boiling
  • 88694 10/21/02 22:17 A. Camsonne Target moved to 15 cm LD2
  • 88690 10/21/02 21:44 A. Camsonne Target Change
  • 88689 10/21/02 21:17 atrig Trigger-Download
  • 88686 10/21/02 21:04 RS,KP,BM,AV Parity Target Boiling Test
  • 88671 10/21/02 18:37 A. Camsonne Target alarm
  • 88657 10/21/02 16:48 R. Michaels end-run halog problems
  • 88654 10/21/02 16:01 Saha Shift Summary
  • 88642 10/21/02 14:33 Eugene Pion rejector
  • 88641 10/21/02 14:23 Saha Magnet Screen for Boiling Studies
  • 88634 10/21/02 13:19 jpchen cryotarget status screen
  • 88633 10/21/02 13:07 jpchen Moved to Lh2-15 cm target position
  • 88632 10/21/02 13:07 moffit Parity Target Boiling Test
  • 88631 10/21/02 12:33 Eugene Pion rejector
  • 88630 10/21/02 12:30 jpchen end controlled access
  • 88629 10/21/02 12:28 R. Michaels replaced ROC2 fastbus busy
  • 88624 10/21/02 11:34 Reitz R-VDC U1 cable rotated
  • 88623 10/21/02 11:08 Suleiman Parity Target Boiling Test
  • 88622 10/21/02 10:34 M Jones carbon current scan
  • 88621 10/21/02 10:05 M Jones 88576
  • 88616 10/21/02 09:18 M Jones found problem in right vdc
  • 88608 10/21/02 08:05 C. Furget an Owl Shift Summary
  • 88600 10/21/02 07:28 atrig Trigger-Download
  • 88599 10/21/02 07:02 atrig Trigger-Download
  • 88598 10/21/02 07:00 atrig Trigger-Download
  • 88597 10/21/02 06:54 M JOnes maybe vally in delta is mapping problem
  • 88596 10/21/02 06:50 H. Ibrahim Harp 5 scan run # 1883
  • 88595 10/21/02 05:01 atrig Trigger-Download
  • 88590 10/21/02 03:50 atrig Trigger-Download
  • 88585 10/21/02 02:55 atrig Trigger-Download
  • 88583 10/21/02 02:14 CF 2580 is Ela_H04 pointing run
  • 88579 10/21/02 01:42 H. Ibrahim Ela_H03 left momentum
  • 88578 10/21/02 01:37 atrig Trigger-Download
  • 88577 10/21/02 01:36 atrig Trigger-Download
  • 88576 10/21/02 00:59 ulmer spec_r_dp spectrum with valley
  • 88575 10/21/02 00:54 ulmer VDC wire spectra
  • 88574 10/21/02 00:36 Eric Shift summary
  • Sunday

  • 88554 10/20/02 14:36 b raue T5 problem?
  • 88549 10/20/02 11:36 jones gain matched left arm gas cerenkov
  • 88548 10/20/02 10:42 jones gain matched right arm A1
  • 88539 10/20/02 08:33 b. raue start of shift
  • 88536 10/20/02 08:01 PU Owl shift summary
  • 88522 10/20/02 04:50 atrig Trigger-Download
  • 88509 10/20/02 02:43 R. Michaels ped run 2566
  • 88508 10/20/02 02:43 R. Michaels DAQ problems and solution
  • Saturday

  • 88463 10/19/02 23:51 b. raue bullseye
  • 88426 10/19/02 22:32 b. raue beam on, wayne
  • 88411 10/19/02 17:48 M JOnes Day shift summary
  • 88408 10/19/02 17:05 epstein beam permit
  • 88402 10/19/02 15:14 H. Ibrahim Target cooled down
  • 88397 10/19/02 14:59 H. Ibrahim Target Controls
  • 88396 10/19/02 14:03 atrig Trigger-Download
  • 88395 10/19/02 14:02 atrig Trigger-Download
  • 88390 10/19/02 12:06 jones rebooted roc2
  • 88387 10/19/02 12:03 jones fixed t2
  • 88386 10/19/02 10:21 M Jones no T2 triggers
  • Friday

  • 88377 10/18/02 11:28 Bob & Bodo trigger & DAQ ready
  • 88375 10/18/02 10:55 atrig Trigger-Download
  • Thursday

  • 88374 10/17/02 15:20 atrig Trigger-Download
  • Wednesday

  • 88371 10/16/02 15:01 R. Michaels experiment switchover
  • 88370 10/16/02 10:06 gilman target skewness
  • 88369 10/16/02 07:14 gilman turning off FPP
  • 88367 10/16/02 07:06 F.Benmokhtar Last E00007 Shift Summary.
  • 88366 10/16/02 07:05 gilman end of experiment
  • 88360 10/16/02 05:08 gilman carbon doors
  • 88358 10/16/02 04:54 gilman kinematics change to 110 cm
  • 88357 10/16/02 04:50 gilman missing wires in fpp front u
  • 88355 10/16/02 04:25 Fatiha Y_Target offset
  • 88353 10/16/02 03:59 gilman kinematics change
  • 88347 10/16/02 02:44 paschke/secr Injector charge asymmetry study
  • 88346 10/16/02 02:44 paschke/secr Injector charge asymmetry, G0 beam off.
  • 88345 10/16/02 02:44 paschke/secr Injector charge asymmetry, G0 beam on
  • 88342 10/16/02 02:41 paschke/secr Study of injector charge asymmetry
  • 88337 10/16/02 01:22 F.Benmokhtar Groups of Wires not responding in the straw chambers.
  • 88329 10/16/02 00:00 Lindgren shift summary
  • Tuesday

  • 88326 10/15/02 23:42 Lindgren production running
  • 88317 10/15/02 22:16 Lindgren Production running
  • 88306 10/15/02 20:50 Lindgren production running
  • 88297 10/15/02 19:10 Lindgren Production running
  • 88291 10/15/02 18:23 Lindgren production running
  • 88287 10/15/02 17:59 Lindgren Production running
  • 88284 10/15/02 17:31 R.A. LIndgre Change targets- production running
  • 88279 10/15/02 16:00 moffit Day Shift Summary
  • 88275 10/15/02 12:59 moffit 30uA Beam Returns
  • 88271 10/15/02 12:14 moffit Dipole is now set correctly
  • 88270 10/15/02 11:57 moffit Re-Setting Left arm momenta
  • 88269 10/15/02 11:48 moffit Some beam returns
  • 88268 10/15/02 09:43 moffit Left-Right momentum set to zero
  • 88266 10/15/02 09:01 moffit Machine Development Time
  • 88265 10/15/02 08:36 gilman move to 70 deg cm
  • 88263 10/15/02 08:22 moffit Back to CW
  • 88261 10/15/02 08:16 moffit Pulsed Beam
  • 88260 10/15/02 08:04 paschke pzt test continue
  • 88258 10/15/02 08:00 F. Benmokhta Shift Summary
  • 88257 10/15/02 07:59 gilman spot for run 1506
  • 88253 10/15/02 07:27 fatiha online results for LD2
  • 88219 10/15/02 02:46 Fatiha y_tg spectrum
  • 88210 10/15/02 01:47 Fatiha FPP Straw Map L.x1.straw and Lx2.straw
  • 88203 10/15/02 00:56 moffit Summary of pzt and ia scans
  • 88194 10/15/02 00:02 vasulk Swing Shift Summay
  • Monday

  • 88186 10/14/02 21:59 vasulk Beam shadow gone
  • 88183 10/14/02 21:04 moffit Shadow in parity daq bcm
  • 88182 10/14/02 21:01 vasulk Beam Spot shadow
  • 88179 10/14/02 19:49 Eugene Moller results
  • 88178 10/14/02 19:44 moffit Begin PZT and IA scans
  • 88175 10/14/02 17:20 olivier 41 uA !
  • 88171 10/14/02 16:04 LeRose Shift Summary
  • 88163 10/14/02 11:01 LeRose Magnet screen
  • 88162 10/14/02 10:46 LeRose Q1 IOC reboot
  • 88161 10/14/02 10:41 gilman CH2 put in
  • 88159 10/14/02 09:49 Lingyan minor changes to FPP eff_front/rear.dat
  • 88154 10/14/02 08:02 F.Benmokhtar Shift Summary
  • 88146 10/14/02 03:51 Fatiha Current Screen
  • 88145 10/14/02 03:43 Fatiha Control screen
  • 88140 10/14/02 00:46 choi Both Q1 IOC problem
  • 88137 10/14/02 00:03 moffit swing shift summary
  • Sunday

  • 88114 10/13/02 16:32 moffit Target moved to LD2
  • 88111 10/13/02 16:05 gilman dummy cell rates
  • 88080 10/13/02 07:59 F.Benmokhtar Shift Summary
  • 88071 10/13/02 05:23 kinney FPP straw efficiency recovery
  • 88061 10/13/02 02:47 kinney FPP straw chamber effic drop
  • 88054 10/13/02 01:11 Jack Argon gas
  • 88053 10/13/02 01:10 F. Benmokhta Low Argon level
  • 88052 10/13/02 00:54 F. Benmokhta Gas sensors alarm
  • 88046 10/13/02 00:00 Punjabi Shift Summary
  • Saturday

  • 88041 10/12/02 22:35 punjabi magnet screeen
  • 88040 10/12/02 22:34 olivier CH2 plug out
  • 88023 10/12/02 17:56 punjabi MGNET SCREEN
  • 87977 10/12/02 08:01 F. Benmokhta Shift Summary
  • 87971 10/12/02 07:07 kinney good FPP scattering event rate
  • 87968 10/12/02 06:52 gilman good FPP scattering event rate
  • 87953 10/12/02 01:58 kinney y_tg LD2, Rad In
  • 87950 10/12/02 01:09 paschke small spike in target vacuum
  • 87948 10/12/02 00:56 Fatiha y_tg for LD2 rad out
  • 87944 10/12/02 00:16 F. Garibaldi swing shift
  • Friday

  • 87915 10/11/02 16:04 bogdanw shift summary
  • 87902 10/11/02 08:03 kinney Owl Shift report 11Oct02
  • 87876 10/11/02 00:01 F.Benmokhtar Shift Summary
  • Thursday

  • 87872 10/10/02 21:53 Fatiha. Coda Rebooted
  • 87870 10/10/02 21:12 Fatiha Spot size.
  • 87869 10/10/02 21:11 jpchen y-tg no big peak with reduced raster size.
  • 87863 10/10/02 19:15 jpchen reduced raster size to 1x1mm
  • 87862 10/10/02 19:03 jpchen target ioc reboot
  • 87860 10/10/02 18:49 jpchen target/beam position and upstream window peak
  • 87856 10/10/02 18:05 olivier target up by 2mm
  • 87852 10/10/02 17:46 jpchen Y-tg with beam shifted down by 1mm
  • 87845 10/10/02 16:02 Higinbotham Shift Summary
  • 87844 10/10/02 16:02 Higinbotham Changed to 3rd Kinematics
  • 87843 10/10/02 16:01 Higinbotham Parasite Data
  • 87829 10/10/02 12:23 Higinbotham tkABU doesn't start
  • 87828 10/10/02 11:59 Higinbotham (0.0,-1.0)
  • 87826 10/10/02 11:57 Higinbotham Pedestal Run
  • 87813 10/10/02 09:55 X. Jiang charge asymmetry in run 1287
  • 87807 10/10/02 08:01 A. Camsonne Owl shift summary
  • 87805 10/10/02 06:11 A. Camsonne Vaccum problem follow up
  • 87804 10/10/02 05:29 A. Camsonne MCC vaccum problem
  • 87798 10/10/02 04:30 A.Camsonne Right arm set at 1.0577 GeV/c
  • 87796 10/10/02 04:07 A. Camsonne Finished quad cycling
  • 87794 10/10/02 03:47 A. Camsonne Right arm cycling quad started
  • 87787 10/10/02 02:45 A. Camsonne Sieve slits in Right Arm
  • 87780 10/10/02 01:44 A. Camsonne Momentum Change to 438.51 MeV/c Right Arm
  • Wednesday

  • 87768 10/09/02 23:58 perdrisa shift log
  • 87756 10/09/02 21:59 R. Michaels updated scaler.map
  • 87753 10/09/02 21:38 Perdrisat L target Y anomaly
  • 87752 10/09/02 21:27 Adaq Hall A general tools screen
  • 87741 10/09/02 19:46 Eugene Beam asymmetry from Moller
  • 87740 10/09/02 19:41 Eugene Moller results
  • 87738 10/09/02 19:34 peter Parity, Moeller & Spectrometer DAQs
  • 87732 10/09/02 16:03 Higinbotham adaql2 not adaql1
  • 87730 10/09/02 15:56 Higinbotham Shift Summary
  • 87729 10/09/02 15:44 Higinbotham Magnet Screen (kin01)
  • 87727 10/09/02 15:16 Higinbotham Target Motion Failure
  • 87725 10/09/02 14:55 holt cryotarget reboot
  • 87702 10/09/02 10:06 R. Michaels BCM D3 missing on L-arm
  • 87699 10/09/02 09:35 Higinbotham adaql1 fixed
  • 87689 10/09/02 08:39 Lindgren Shift Summary
  • 87688 10/09/02 08:39 gilman T3 rates
  • 87687 10/09/02 08:26 gilman reboot
  • 87686 10/09/02 08:05 Lindgren Shift Summary
  • 87684 10/09/02 07:58 R. Michaels adaql2 stuck 'ls' problem
  • 87683 10/09/02 07:41 R. Michaels halogscript stuck
  • 87682 10/09/02 04:10 slifer CODA end of run delay
  • 87681 10/09/02 03:59 slifer left arm ytg
  • 87680 10/09/02 03:41 Lindgren Mid-Shift Summary
  • 87679 10/09/02 03:14 Adaq ytarg spectra
  • 87678 10/09/02 02:36 slifer CODA
  • 87677 10/09/02 02:33 slifer Left arm Scintillator TDCs
  • 87676 10/09/02 01:28 Lindgren Production running, computer crash
  • 87675 10/09/02 00:04 bogdanw shift summary
  • Tuesday

  • 87667 10/08/02 22:22 Issam Qattan Funny 9 foils Optics target
  • 87655 10/08/02 20:31 olivier optics target fixed
  • 87650 10/08/02 20:15 Lhuillier Bkg in Compton chicane
  • 87644 10/08/02 18:36 deleo Fields
  • 87641 10/08/02 18:06 R. Michaels event type 140's had problems
  • 87640 10/08/02 17:20 R. Michaels charge asymmetry test
  • 87639 10/08/02 17:19 bogdanw right HV
  • 87638 10/08/02 17:18 bogdanw left HV
  • 87635 10/08/02 16:08 Eugene Moller results
  • 87634 10/08/02 16:07 Eugene Moller measurement
  • 87633 10/08/02 16:06 Ole Hansen Shift Summary
  • 87631 10/08/02 10:00 R. Michaels Re: DAQ problems (TS0)
  • 87630 10/08/02 09:51 gayou better screen grab
  • 87629 10/08/02 09:49 gilman rates
  • 87628 10/08/02 09:38 gayou figure of y-tgt-left
  • 87627 10/08/02 09:24 olivier y_tg shows radiator
  • 87625 10/08/02 09:08 Higinbotham Magnet Screen
  • 87623 10/08/02 08:55 Higinbotham ps3=ps5=1
  • 87622 10/08/02 08:54 gilman spot offset?
  • 87621 10/08/02 08:51 Ole Hansen Fast Feedback not operational
  • 87619 10/08/02 08:49 Higinbotham Injector Optics Correction
  • 87614 10/08/02 08:19 Ole Hansen Fast Feedback
  • 87613 10/08/02 08:08 Lindgren shift summary
  • 87610 10/08/02 07:30 solvigno coulomb sum rule test
  • 87605 10/08/02 06:17 Lindgren Computer crash
  • 87604 10/08/02 06:10 solvigno TS0 is back
  • 87602 10/08/02 05:53 solvigno TS0 crashed and have it back
  • 87592 10/08/02 03:42 A. Camsonne Target Moved to D2 / Radiator
  • 87591 10/08/02 03:34 Lindgren beam,target
  • 87587 10/08/02 02:18 A. Camsonne Target movement fixed
  • 87586 10/08/02 00:41 A. Camsonne Target movement problem
  • 87585 10/08/02 00:24 olivier Wien angle changed
  • 87583 10/08/02 00:00 olivier wien angle not changed
  • Monday

  • 87582 10/07/02 21:57 R. Michaels DAQ checks, start cosmics run 1215
  • 87576 10/07/02 21:00 gilman resetting right arm rocs
  • 87575 10/07/02 20:16 K. McCormick Left Arm Dipole Quench
  • 87574 10/07/02 17:23 Jack R-Arm Shower HV
  • 87573 10/07/02 16:24 olivier HV set for left aerogel
  • 87572 10/07/02 16:16 olivier right A2 HV screen (2)
  • 87571 10/07/02 16:15 olivier right A2 HV screen (1)
  • 87570 10/07/02 16:11 olivier right arm HV
  • 87569 10/07/02 16:08 olivier move spectrometers
  • 87568 10/07/02 16:06 jpchen Day shift summary
  • 87566 10/07/02 15:29 jpchen optics target position
  • 87565 10/07/02 14:53 K. McCormick Moved test wire chamber to 105 degrees
  • 87564 10/07/02 14:53 jpchen charts etc.
  • 87563 10/07/02 14:40 olivier fpp spectrum understood (and normal)
  • 87562 10/07/02 14:12 jpchen cryotarget cool-down completed
  • 87561 10/07/02 12:02 R. Michaels BPM pedestal run
  • 87560 10/07/02 11:55 R. Michaels 'spot' works for fastbus data now
  • 87559 10/07/02 10:38 Jack R-Arm VDC cable reseated
  • 87556 10/07/02 09:14 olivier fpp carbon doors
  • 87555 10/07/02 08:35 jpchen cryotarget chart diplay of vacuum and current not working
  • 87554 10/07/02 08:33 jpchen Hall A in restricted access now for target switch
  • 87551 10/07/02 07:13 Lindgren shift summary
  • 87550 10/07/02 07:07 jpchen 15K supply valve kept open.
  • 87549 10/07/02 07:06 jpchen Cryotarget chart scale.
  • 87548 10/07/02 07:04 jpchen cryotarget warm up
  • 87526 10/07/02 00:42 olivier fpp spectrum
  • 87521 10/07/02 00:02 A. Deur Shift Summary
  • Sunday

  • 87518 10/06/02 23:29 A. Deur The Q1s tripped again
  • 87517 10/06/02 23:02 Issam Qattan Changing header file for kin#4
  • 87501 10/06/02 22:03 A. Deur ep4-3
  • 87500 10/06/02 21:59 A. Deur Setup for kinematics ep4-4
  • 87498 10/06/02 21:58 A. Deur Done with kinematics ep4-2.
  • 87497 10/06/02 21:51 A. Deur FPP carbon doors
  • 87495 10/06/02 20:38 A. Deur C12 optic target
  • 87485 10/06/02 16:55 X. Jiang helicity in run 1173
  • 87484 10/06/02 16:51 X. Jiang changed helicity algorithm, run 1170
  • 87480 10/06/02 16:22 gilman wien angle
  • 87479 10/06/02 16:01 Ransome shift summary
  • 87477 10/06/02 15:56 gilman beam tweaking
  • 87474 10/06/02 15:07 Ransome LH2 target position question
  • 87471 10/06/02 14:48 R. Michaels helicity disaster (NOT)
  • 87468 10/06/02 13:56 R. Michaels helicity errors reported, run 1173
  • 87465 10/06/02 12:48 olivier tc_cor cut
  • 87463 10/06/02 12:23 Ransome Q2-L trip
  • 87458 10/06/02 11:59 R. Michaels some shifted VDC data L-arm
  • 87456 10/06/02 10:49 olivier S2-L-4 fixed
  • 87455 10/06/02 10:42 R. Michaels DAQ deadtime in UNBUFF
  • 87454 10/06/02 10:40 Ransome slight increase in current
  • 87453 10/06/02 10:36 ransome change of kinematics
  • 87451 10/06/02 10:16 R. Michaels Switch to kine ep4-3, test runs 1167-68
  • 87448 10/06/02 10:03 R. Michaels re: synch problems
  • 87446 10/06/02 09:37 R. Michaels detector problems - follow up
  • 87445 10/06/02 09:18 R. Michaels run 1165 ended in crash
  • 87443 10/06/02 09:07 Ransome beam position
  • 87442 10/06/02 09:06 Ransome entries in log book
  • 87441 10/06/02 08:57 Ransome beam position
  • 87437 10/06/02 08:01 Dutta 10/06/02 Owl Shift Summary
  • 87427 10/06/02 04:54 dutta configuration / kinematics change
  • 87426 10/06/02 04:40 moffit Y-Target Left and Right Spec
  • 87424 10/06/02 03:52 gilman gilman
  • 87414 10/06/02 01:57 ddutta R Arm Q1 trips
  • 87412 10/06/02 01:32 moffit TDC double shelf
  • 87411 10/06/02 01:13 moffit End of run#1153 failed
  • 87407 10/06/02 00:57 moffit ROC and TS crash
  • 87406 10/06/02 00:29 A. Deur Shift summary
  • Saturday

  • 87402 10/05/02 23:29 X. Jiang helicity flag in run 1150
  • 87401 10/05/02 23:25 X. Jiang run 1150, helicity OK
  • 87396 10/05/02 22:19 A. Deur Drop the linearity test.
  • 87395 10/05/02 22:07 olivier optics target unavailable
  • 87390 10/05/02 21:26 A. Deur carbon optic target
  • 87389 10/05/02 20:35 A. Deur bcm was not working
  • 87388 10/05/02 20:03 olivier ep ion chamber masked for beam tuning
  • 87387 10/05/02 18:09 A. Deur Right collimator out
  • 87386 10/05/02 17:56 x. jiang push CH2 analyser in
  • 87385 10/05/02 17:25 x. jiang carbon doors are driven in
  • 87381 10/05/02 17:06 x. jiang FPP low voltage screen
  • 87380 10/05/02 17:05 X. Jiang FPP high voltage and low voltage screen
  • 87375 10/05/02 15:55 Ransome shift summary
  • 87371 10/05/02 15:13 R. Michaels turned off event limit
  • 87361 10/05/02 10:51 R. Michaels detector problems (small)
  • 87359 10/05/02 10:21 X. jiang checked detectors, run 1137
  • 87356 10/05/02 10:11 R. Michaels miscellaneous / HOWTO for BPM pedestals
  • 87354 10/05/02 09:05 Ransome Q1 trip, L and R
  • 87351 10/05/02 08:02 moffit shift summary
  • 87348 10/05/02 07:35 moffit target moved at MCC request
  • 87344 10/05/02 03:17 moffit pulsed beam
  • 87343 10/05/02 03:03 moffit Left and Right Q1s reset
  • 87342 10/05/02 02:34 moffit controlled access and cryo
  • 87341 10/05/02 02:14 gilman cryotarget status
  • 87340 10/05/02 01:29 moffit controlled access
  • 87339 10/05/02 01:20 gilman cryotarget problem?
  • 87338 10/05/02 00:13 A. Deur Shift Summary
  • Friday

  • 87336 10/04/02 23:57 gilman target moved at MCC request
  • 87335 10/04/02 22:51 A. Deur update on the experiment
  • 87334 10/04/02 21:51 A. Deur ion chamber trip
  • 87333 10/04/02 21:41 ransome shift update, Q1 reset
  • 87332 10/04/02 20:59 Reitz raster, bpm and espace
  • 87331 10/04/02 20:57 Olivier Access for Q1R
  • 87330 10/04/02 20:20 Ransome update on shift
  • 87329 10/04/02 20:19 A. Deur Trouble locking the NMR for left HRS
  • 87328 10/04/02 20:14 Reitz Harp Scans Failed
  • 87321 10/04/02 19:41 Ransome update on shift
  • 87316 10/04/02 16:09 jpchen Day Shift Summary
  • 87309 10/04/02 13:05 gilman FPP chamber 2 fuse blown
  • 87305 10/04/02 09:52 jpchen Hall A in restricted access
  • 87303 10/04/02 08:01 Eric Voutier Shift Summary
  • 87296 10/04/02 00:29 A. Deur Target motion failure
  • 87294 10/04/02 00:08 Olivier Night program
  • 87293 10/04/02 00:06 A. Deur Shift Summary
  • Thursday

    
    

    Tuesday

  • 87281 10/01/02 10:25 atrig Trigger-Download
  • 87280 10/01/02 10:23 atrig Trigger-Download